OpenStudy (daryan):

ln(x+5)=ln(x+8) - ln(x+4) solve for x please

OpenStudy (solomonzelman):

\(\Large\color{blue}{ \bf ln(x+5)=ln(x+8) - ln(x+4) }\) \(\Large\color{blue}{ \bf ln(x+5)+ ln(x+4)=ln(x+8) }\) \(\Large\color{blue}{ \bf ln[~~(x+5)(x+4)~~]=ln(x+8) }\) \(\Large\color{blue}{ \bf ln[~~x^2+9x+20~~]=ln(x+8) }\) \(\Large\color{red}{ \bf x^2+9x+20=x+8 }\)

OpenStudy (solomonzelman):

you got the rest, go for it .... :)

OpenStudy (daryan):

My answer choices are a) -10 and -2, b) -2, c) -2 and 8, d) -10 and 2

OpenStudy (jdoe0001):

ahemm solve for "x"

OpenStudy (solomonzelman):

Starting from what I wrote in red. 1) subtract x from both sides 2) subtract 8 from both sides 3) solve by factoring/completing the square (same thing) you got this...

OpenStudy (daryan):

I get -2 and -6 but that is not in my possible choices